# MPMA: Preference Manipulation Attack Against Model Context Protocol

### v1

@@ Appendix D Ethical Consideration
Appendix E Limitations
Although the proposed MPMA attack demonstrates notable effectiveness, several limitations must be acknowledged to fully understand its practical applicability. First, the success of the attack heavily depends on the internal selection mechanism of the target LLM. However, this mechanism is essentially a black box, making it exhibit limited interpretability. Specifically, experiment results reveal significant differences in ASR across different models, such as Gemini[53] and GPT-4o[43]. While the underlying cause remains unclear, it is hypothesized that these variations may stem from differences in internal defense strategies or heightened sensitivity to linguistic manipulation. Second, the scope of evaluation is limited. Although this study includes five mainstream LLMs and eight commonly used MCP servers, the lack of automation in the MCP evaluation process results in slow experiment throughput. Broader experiment coverage is needed in future work to assess the more universal applicability of MPMA. Third, the reproducibility of the results is affected by the inherent randomness in LLM outputs. Due to the stochastic nature of LLM responses, identical inputs may yield different selection behaviors at different times, thereby making the complete reproduction of this study particularly complex.
Appendix F Broader Impacts
The MPMA attack framework proposed in this study is not only technically insightful but also introduces a range of societal risks. From a negative perspective, MPMA poses a tangible threat to the security of open LLM agents and MCP ecosystems. By embedding direct or invisible manipulations into MCP server descriptions, an attacker can bias the modelβs tool selection process to favor their own services and thereby gain economic benefits. Such manipulation has the potential to exacerbate existing social inequalities. If malicious tools are deployed at scale, the attack could result in monopolization of user queries by the attacker, undermining the fairness and diversity of the ecosystem. From a positive standpoint, the public disclosure of this research provides valuable insights for identifying and addressing security vulnerabilities in MCP and LLM agent systems. By presenting the attack methodology and evaluation mechanisms, the study aims to foster the development of corresponding defense strategies within the community. To mitigate the risk of misuse, access to the code and dataset has been restricted to research purposes only, accompanied by clear warnings regarding potential abuse. Furthermore, the study recommends the incorporation of trusted labeling mechanisms in future MCP infrastructures to enhance platform security and bolster user trust.

### v2

@@ Appendix D Ethical Consideration
Appendix E Broader Impacts
The MPMA attack framework proposed in this study is not only technically insightful but also introduces a range of societal risks. From a negative perspective, MPMA poses a tangible threat to the security of open LLM agents and MCP ecosystems. By embedding direct or invisible manipulations into MCP server descriptions, an attacker can bias the model’s tool selection process to favor their own services and thereby gain economic benefits. Such manipulation has the potential to exacerbate existing social inequalities. If malicious tools are deployed at scale, the attack could result in monopolization of user queries by the attacker, undermining the fairness and diversity of the ecosystem. From a positive standpoint, the public disclosure of this research provides valuable insights for identifying and addressing security vulnerabilities in MCP and LLM agent systems. By presenting the attack methodology and evaluation mechanisms, the study aims to foster the development of corresponding defense strategies within the community. To mitigate the risk of misuse, access to the code and dataset has been restricted to research purposes only, accompanied by clear warnings regarding potential abuse. Furthermore, the study recommends the incorporation of trusted labeling mechanisms in future MCP infrastructures to enhance platform security and bolster user trust.